# Orbital Altermagnetism

### v1

@@ Abstract
In this Letter, we introduce the concept of orbital altermagnetism—an altermagnetic order originating from staggered local orbital magnetic moments, namely, a magnetic ordered phase of pure orbital degrees of freedom. Using a minimal tight-binding model with spontaneously ordered loop currents, we demonstrate the emergence of a characteristic real-space pattern of anti-parallel orbital magnetic moments and a distinct feature of d-wave-like orbital-momentum locking in momentum space. Importantly, we show that orbital altermagnetism can exist independent of spin configurations in realistic materials. First-principles calculations reveal a definite manifestation of orbital altermagnetism in in-plane ferromagnets CuBr2 and VS2 with staggered out-of-plane orbital moments distinct from the uniform in-plane spin distributions. Furthermore, we find that conventional spin altermagnets can be accompanied with a hidden orbital altermagnetism, as exemplified by monolayer MoO and CrO, and we uncover a giant nonlinear current-induced orbital magnetization that far exceeds the spin contribution, which may serve as an experimental signature of the orbital altermagnetism.
@@
Figure 2: (a) Top and side views of the monolayer CuBr2. Spin magnetic moments localized on Cu atoms are annotated with red arrows. (b) Band structures of CuBr2, with the color denoting the magnitude of $L_{z}$ and $S_{y}$ for each band. The inset displays the first Brillouin zone and high-symmetry points, with $S_{1}=(0.5,0.5)$ and $S_{2}=(0.5,-0.5)$. (c) Real-space distribution of orbital magnetic moments in a $10\times 10$ lattice under open boundary conditions at $\mu=0$ eV. Red (blue) denotes positive (negative) $M_{z}({\bm{r}})$. (d) Momentum-space distribution of $L_{z}({\bm{k}})$ for the lowest conduction band (the 43rd band) across the Brillouin zone.
@@
A prototypical site-type example is CuBr2 [80]. In its ground state [Fig. 2(a)], Cu spins align ferromagnetically along $y$, giving the magnetic point group $2/m$. The two Cu atoms are related by either a glide mirror $\widetilde{m}_{y}=\{m_{y}|\tfrac{1}{2},\tfrac{1}{2},0\}$ or a glide rotation $\widetilde{C}_{2y}=\{C_{2y}|\tfrac{1}{2},\tfrac{1}{2},0\}$, both of which reverse $M_{z}$. Our first-principles calculations with SOC indeed reveal opposite-sign $M_{z}({\bm{r}})$ on two Cu atoms in real space, as shown in Fig. 2(c). Furthermore, the band structure along the diagonal and antidiagonal $\bm{k}$}-paths ($S_{1}$-$\Gamma$-$S_{2}$) exhibits opposite $L_{z}$ polarization with finite amplitude ($\sim 0.1\hbar$), while maintaining persistent spin $S_{y}$ polarization [Fig. 2(b)]. The $\bm{k}$}-space distribution of $L_{z}({\bm{k}})$ further manifests a defining characteristic $d$-wave pattern [see Fig. 2(d)]. Rotating the spins to the $x$ axis changes the magnetic group to $2^{\prime}/m^{\prime}$, where the $\widetilde{m}_{y}\mathcal{T}$ symmetry forces identical $M_{z}$ on both Cu atoms, thereby suppressing orbital altermagnetism. Consequently, the presence or absence of orbital altermagnetism for different spin alignments produces distinct nonlinear orbital Edelstein effects to be used as an experimental signature of orbital altermagnetism (see SM footnotemark: ).
Figure 3: (a) Top and side views of the monolayer MoO lattice. Green arrows indicate the spin magnetic moments on Mo atoms. (b) Spin-($S_{z}$) and Orbital- ($L_{z}$) resolved band structures, where nearly degenerated band crossing points at $\mu=0.133$ eV are marked as $W_{1}$ and $W_{2}$. The inset displays the Brillouin zone with high-symmetry points. (c) Momentum-space distribution of $S_{z}({\bm{k}})$ (top panel) and $L_{z}({\bm{k}})$ (bottom panel) for the highest valence band across the Brillouin zone, with $k_{x}$ and $k_{y}$ in reduced coordinates. (d) Real-space distribution of orbital magnetic moments in a $10\times 10$ lattice under open boundary conditions at $\mu$=0 eV. Red (blue) denotes positive (negative) $M_{z}({\bm{r}})$. (e) Nonlinear responses $\chi^{X}_{zyy}$ ($X=L$ or $S$) as a function of chemical potential $\mu$ from first-principles calculations. (f) Momentum-space distribution of $\chi_{zyy}({\bm{k}})$ across the Brillouin zone at $\mu=0.133$ eV.
@@
Orbital altermagnetism in conventional spin altermagnets.—We next illustrate the coexistence of spin and orbital altermagnetism using monolayer MoO as a representative example. The lattice structure [Fig. 3(a)] belongs to the $4^{\prime}/mm^{\prime}m$ magnetic point group. Its band structure exhibits clear momentum-dependent $S_{z}$ and $L_{z}$ polarizations, indicating simultaneous spin and orbital altermagnetic order [Fig. 3(b)]. The alternating momentum-space distribution of $L_{z}({\bm{k}})$ for the highest valence band [Fig. 3(c)] and the staggered real-space arrangement of orbital magnetic moments [Fig. 3(d)] confirm a $d$-wave-type orbital altermagnetism. Similar behavior is also found in monolayer CrO[82], with results presented in SM footnotemark: .
Like its spin counterpart, orbital altermagnetism contributes to a variety of magnetic phenomena, including the magneto-optical Kerr effect [83, 84], X-ray magnetic dichroism [85, 86], piezomagnetism [4, 22], and nonlinear Edelstein effect [87, 88, 89]. For instance, in the nonlinear Edelstein effect, a nonlinear current-induced spin or orbital magnetization can be expressed as $\delta X_{a}=\chi^{X}_{abc}E_{b}E_{c}$, where $a,b,c\in\{x,y,z\}$, $E_{a}$ is the electric field, and $X=S\,\mathrm{or}\,L$ denote the spin or orbital contributions, respectively. The nonlinear response tensor $\chi^{X}_{abc}$ is governed by band-geometric quantities such as the quantum metric, Berry curvature, and the matrix elements of spin or orbital angular momentum (see SM for details footnotemark: ). In monolayer MoO, the $C_{4z}\mathcal{T}$ symmetry of the altermagnetic order enforces a $\pi$-periodic angular dependence, $\chi^{X}_{zy^{\prime}y^{\prime}}(\theta)=\chi^{X}_{zyy}(0)\cos 2\theta$.
@@
Overall, our findings identify orbital altermagnetism as a broad and robust class of magnetic order—one that is distinct from, yet intertwined with, spin altermagnetism—opening a pathway to novel symmetry-driven magnetotransport phenomena and electrically tunable orbital-based spintronic devices. Experimentally, the significant orbital magnetization establishes giant current-induced orbital polarization as a transport signature of orbital altermagnetism and suggests that orbital degrees of freedom may play a dominant role in certain magnetic responses. In addition to conventional magnetic materials studied here, the orbital altermagnetism is also expected to be achieved from the orbital antiferromagnetism in correlated systems, such as cuprates and cerium hexaboride CeB6, via various symmetry-engineering approaches such as strain[96], electric field [97, 98, 99], controllable stacking bilayers [100, 101, 102, 103], ferroelectric substrate [104], and Janusization [99, 105].
@@ Appendix A Formula of Orbital Magnetization in Reciprocal and Real Spaces
In two dimensions, $M_{n{\bm{k}}}$ is along the out-of-plane direction and is directly related to the orbital angular momentum via
@@ Appendix B Magnetic octupole and Magnetic Toroidal quadrupole
For the minimal model defined in Eq. (1), we reveal that $T_{xy}\neq 0$ and $M_{z}^{\beta}\neq 0$ (with $O_{z}^{\beta}=\frac{\sqrt{15}}{2}z(x^{2}-y^{2}))$, consistent with the allowed multipoles of the corresponding magnetic point group [110]. Particularly, the nonzero magnetic octupole $M_{z}^{\beta}$ can serve as a secondary order parameter in the Landau theory of the $d$-wave altermagnetism [68]. The nonzero $T_{xy}$ can also be directly traced back to the imaginary hopping terms in the tight-binding model. By evaluating the symmetry-adapted multipole basis (SAMB) for the bond cluster [111], we find that in the D4h point group, only the irreducible representation B2g yields a non-zero toroidal quadrupole, which precisely corresponds to $T_{xy}$.
@@ Appendix C Intrinsic nonlinear current-induced spin and orbital magnetization
In this section, we briefly introduce the formalism of the intrinsic nonlinear Edelstein effect. For a 2D system with either $C_{2z}$ or $\mathcal{P}$ symmetry, the linear spin and orbital responses are forbidden by symmetry. In such cases, the leading contribution arises at the second order:
@@
where $X_{a}$ denotes the $a$-th component of either the orbital angular momentum $L_{a}$ or the spin angular momentum $S_{a}$. The nonlinear response tensor $\chi_{abc}$ can be decomposed according to its parity under time-reversal symmetry [87, 88, 89]:
$\begin{split}\chi_{abc}^{\mathrm{even}}&=\tau\sum_{n}\int[d{\bm{k}}]f_{0}\partial_{c}\gamma^{n}_{ab},\\ \chi_{abc}^{\mathrm{odd}}&=\frac{1}{2}\int[d{\bm{k}}]\sum_{n}\Bigg[\Lambda_{abc,n}^{X}f_{0}-(X_{a,n}G_{bc,n}-v_{b,n}\mathscr{G}_{ac,n}-v_{c,n}\mathscr{G}_{ab,n})f^{\prime}_{0}\Bigg].\end{split}$ (10)
@@
The Zeeman-like field $\bm{h}$ couples to the spin or orbital degree of freedom in the form of $\bm{\hat{s}}\cdot\bm{h}$ or $\bm{\hat{L}}\cdot\bm{h}$. Therefore, in the presence of an external field, one can define the Berry connection and the related quantities in the $h$-space [87]. For example, the $h$-space Berry connection is defined as $\mathfrak{A}_{i}=\langle u_{n}({\bm{k}})|i\partial_{h_{i}}|u_{n}({\bm{k}})\rangle$. The Berry connection polarizability (BCP) and the $h$-space BCP are defined as
$\begin{split}G_{ab,n}&=2\textnormal{Re}\sum_{m\neq n}\frac{v_{a,nm}v_{b,mn}}{(\epsilon_{n}-\epsilon_{m})^{3}},\\ \mathscr{G}_{ab,n}&=-2\textnormal{Re}\sum_{m\neq n}\frac{s_{a,nm}v_{b,mn}}{(\epsilon_{n}-\epsilon_{m})^{3}},\end{split}$ (11)
where the numerator involves the interband matrix elements of spin and velocity operators, $s_{a,nm}$ and $v_{a,nm}$. The anomalous spin polarizability (ASP) is given by
$\gamma^{n}_{ab}({\bm{k}})=2\textnormal{Im}\sum_{m\neq n}\frac{s_{a,nm}v_{b,mn}}{(\epsilon_{n}-\epsilon_{m})^{2}}.$ (12)
@@
$\begin{split}\langle u_{n}|\hat{\bm{L}}|u_{m}\rangle&=i\frac{e\hbar^{2}}{4g_{L}\mu_{B}}\sum_{q\neq n,m}(\frac{1}{\epsilon_{q}-\epsilon_{n}}+\frac{1}{\epsilon_{q}-\epsilon_{m}})\cdot\langle u_{n}|\hat{\bm{v}}|u_{q}\rangle\times\langle u_{q}|\hat{\bm{v}}|u_{m}\rangle.\end{split}$ (14)

### v2

@@ Abstract
In this Letter, we introduce the concept of orbital altermagnetism—an altermagnetic order originating from staggered local orbital magnetic moments, namely, a magnetic ordered phase of pure orbital degrees of freedom [52, 53, 54]. Using a minimal tight-binding model with spontaneously ordered loop currents, we demonstrate the emergence of a characteristic real-space pattern of anti-parallel orbital magnetic moments and a distinct feature of d-wave-like orbital-momentum locking in momentum space. Importantly, we show that orbital altermagnetism can exist independent of spin configurations in realistic materials. First-principles calculations reveal a definite manifestation of orbital altermagnetism in in-plane ferromagnets CuBr2 and VS2 with staggered out-of-plane orbital moments distinct from the uniform in-plane spin distributions. Furthermore, we find that conventional spin altermagnets can be accompanied with a hidden orbital altermagnetism, as exemplified by monolayer MoO and CrO, and we uncover a giant nonlinear current-induced orbital magnetization that far exceeds the spin contribution, which may serve as an experimental signature of the orbital altermagnetism.
@@
Figure 2: (a) Top and side views of the monolayer CuBr2. Spin magnetic moments localized on Cu atoms are annotated with red arrows. (b,c) Band structures of CuBr2, with the color denoting the magnitude of $L_{z}$ (b) and $S_{y}$ (c) for each band. The inset displays the first Brillouin zone and high-symmetry points, with $S_{1}=(0.5,0.5)$ and $S_{2}=(0.5,-0.5)$. (d) Real-space distribution of orbital magnetic moments in a $10\times 10$ lattice under open boundary conditions at $\mu=0$ eV. Red (blue) denotes positive (negative) $M_{z}({\bm{r}})$. (e) Momentum-space distribution of $L_{z}({\bm{k}})$ for the lowest conduction band (the 43rd band) across the Brillouin zone.
@@
A prototypical site-type example is CuBr2 [83]. We consider an in-plane ferromagnetic configuration [Fig. 2(a)], where the Cu spins align ferromagnetically along $y$, giving the magnetic point group $2/m$. The two Cu atoms are related by either a glide mirror $\widetilde{m}_{y}=\{m_{y}|\tfrac{1}{2},\tfrac{1}{2},0\}$ or a glide rotation $\widetilde{C}_{2y}=\{C_{2y}|\tfrac{1}{2},\tfrac{1}{2},0\}$, both of which reverse $M_{z}$. Our first-principles calculations with SOC indeed reveal opposite-sign $M_{z}({\bm{r}})$ on two Cu atoms in real space, as shown in Fig. 2(d). Furthermore, the band structure along the diagonal and antidiagonal $\bm{k}$}-paths ($S_{1}$-$\Gamma$-$S_{2}$) exhibits opposite $L_{z}$ polarization with finite amplitude ($\sim 0.1\hbar$) [Fig. 2(b)], while maintaining persistent spin $S_{y}$ polarization [Fig. 2(c)]. The $\bm{k}$}-space distribution of $L_{z}({\bm{k}})$ further manifests a defining characteristic $d$-wave pattern [see Fig. 2(e)]. Rotating the spins to the $x$ axis changes the magnetic group to $2^{\prime}/m^{\prime}$, where the $\widetilde{m}_{y}\mathcal{T}$ symmetry forces identical $M_{z}$ on both Cu atoms, thereby suppressing orbital altermagnetism. Consequently, the presence or absence of orbital altermagnetism for different spin alignments produces distinct nonlinear orbital Edelstein effects to be used as an experimental signature of orbital altermagnetism (see SM footnotemark: ).
Figure 3: (a) Top and side views of the monolayer MoO lattice. Green arrows indicate the spin magnetic moments on Mo atoms. (b) Spin-($S_{z}$) and (d) Orbital- ($L_{z}$) resolved band structures, where nearly degenerated band crossing points at $\mu=0.133$ eV are marked as $W_{1}$ and $W_{2}$. The inset displays the Brillouin zone with high-symmetry points. (c,e) Momentum-space distribution of $S_{z}({\bm{k}})$ (c) and $L_{z}({\bm{k}})$ (e) for the highest valence band across the Brillouin zone, with $k_{x}$ and $k_{y}$ in reduced coordinates. Note that the color maps of $L_{z}({\bm{k}})$ are presented in logarithmic scale for better visualization of its $\bm{{\bm{k}}}$-space variation. (f) Real-space distribution of orbital magnetic moments in a $10\times 10$ lattice under open boundary conditions at $\mu$=0 eV. Red (blue) denotes positive (negative) $M_{z}({\bm{r}})$. (g) Nonlinear responses $\chi^{X}_{zyy}$ ($X=L$ or $S$) as a function of chemical potential $\mu$ from first-principles calculations. (h) Momentum-space distribution of $\chi_{zyy}({\bm{k}})$ across the Brillouin zone at $\mu=0.133$ eV.
@@
Orbital altermagnetism in conventional spin altermagnets.—We next illustrate the coexistence of spin and orbital altermagnetism using monolayer MoO as a representative example. The lattice structure [Fig. 3(a)] belongs to the $4^{\prime}/mm^{\prime}m$ magnetic point group. Its band structure exhibits clear momentum-dependent $S_{z}$ and $L_{z}$ polarizations, indicating simultaneous spin and orbital altermagnetic order [Fig. 3(b,d)]. The alternating momentum-space distribution of $L_{z}({\bm{k}})$ for the highest valence band [Fig. 3(c,e)] and the staggered real-space arrangement of orbital magnetic moments [Fig. 3(f)] confirm a $d$-wave-type orbital altermagnetism. Similar behavior is also found in monolayer CrO[85], with results presented in SM footnotemark: .
Like its spin counterpart, orbital altermagnetism contributes to a variety of magnetic phenomena, including the magneto-optical Kerr effect [86, 87], X-ray magnetic dichroism [88, 89], piezomagnetism [4, 22], and nonlinear Edelstein effect [90, 91, 92]. For instance, in the nonlinear Edelstein effect, a nonlinear current-induced spin or orbital magnetization can be expressed as $\delta X_{a}=\chi^{X}_{abc}E_{b}E_{c}$, where $a,b,c\in\{x,y,z\}$, $E_{a}$ is the electric field, and $X=S\,\mathrm{or}\,L$ denote the spin or orbital contributions, respectively. The nonlinear response tensor $\chi^{X}_{abc}$ is governed by band-geometric quantities such as the quantum metric, Berry curvature, and the matrix elements of spin or orbital angular momentum (see Sec. III SM for details footnotemark: ). In monolayer MoO, the combined symmetries $C_{4z}\mathcal{T}$ and $m_{x}\mathcal{T}$ strongly constrain the second-order response tensor $\chi^{X}_{zij}$. For the out-of-plane response ($z$ index), the time-reversal-even part is symmetry-forbidden, leaving only the intrinsic (time-reversal-odd) contribution. In this case, $m_{x}\mathcal{T}$ enforces $\chi^{X}_{zxy}=0$, while $C_{4z}\mathcal{T}$ imposes $\chi^{X}_{zxx}=-\chi^{X}_{zyy}$. Consequently, for an in-plane electric field $\mathbf{E}=E(\sin\theta,\cos\theta)$, the induced out-of-plane polarization varies as $\delta X_{z}\propto E^{2}\cos 2\theta$, leading to a characteristic $\pi$-periodic angular dependence.
@@
Overall, our findings identify orbital altermagnetism as a broad and robust class of magnetic order—one that is distinct from, yet intertwined with, spin altermagnetism—opening a pathway to novel symmetry-driven magnetotransport phenomena and electrically tunable orbital-based spintronic devices. Experimentally, the significant orbital magnetization establishes giant current-induced orbital polarization as a transport signature of orbital altermagnetism and suggests that orbital degrees of freedom may play a dominant role in certain magnetic responses. We expect our theoretical proposal will stimulate immediate experimental interests, since recent experimental developments have made the detection of the real-space orbital magnetization patterns and altermagnetic orbital texture in momentum space possible, such as circular dichroism combined with soft x-ray angle-resolved photoemission spectroscopy [99], and nitrogen-vacancy (NV)-center magnetometry [100] (see Sec. VII in SM footnotemark: for detailed discussion). In addition to conventional magnetic materials studied here, the orbital altermagnetism is also expected to be achieved from the orbital antiferromagnetism in correlated systems, such as cuprates and cerium hexaboride CeB6, via various symmetry-engineering approaches such as strain[101], electric field [102, 103, 104], controllable stacking bilayers [105, 106, 107, 108], ferroelectric substrate [109], and Janusization [104, 110].
@@ Acknowledgements.
Data availability—The data that support the findings of this article are openly available [111].
@@ Appendix A Formula of Orbital Magnetization in Reciprocal and Real Spaces
More specifically,
$M_{n{\bm{k}}}=-\frac{e}{2\hbar c}\,\mathrm{Im}\!\sum_{m\neq n}\frac{\langle u_{n{\bm{k}}}|\hat{\mathbf{v}}|u_{m{\bm{k}}}\rangle\times\langle u_{m{\bm{k}}}|\hat{\mathbf{v}}|u_{n{\bm{k}}}\rangle}{\varepsilon_{n{\bm{k}}}-\varepsilon_{m{\bm{k}}}},$ (4)
where $\hat{\mathbf{v}}=\frac{1}{\hbar}\nabla_{{\bm{k}}}H({\bm{k}})$ is the velocity operator. From this expression, the orbital magnetic moment is expected to increase significantly in regions where the bands are nearly degenerate. In two dimensions, $M_{n{\bm{k}}}$ is along the out-of-plane direction and is directly related to the orbital angular momentum via
@@ Appendix B Magnetic octupole and Magnetic Toroidal quadrupole
For the minimal model defined in Eq. (1) in the main text, we reveal that $T_{xy}\neq 0$ and $M_{z}^{\beta}\neq 0$ (with $O_{z}^{\beta}=\frac{\sqrt{15}}{2}z(x^{2}-y^{2}))$, consistent with the allowed multipoles of the corresponding magnetic point group [116]. Particularly, the nonzero magnetic octupole $M_{z}^{\beta}$ can serve as a secondary order parameter in the Landau theory of the $d$-wave altermagnetism [71]. The nonzero $T_{xy}$ can also be directly traced back to the imaginary hopping terms in the tight-binding model. By evaluating the symmetry-adapted multipole basis (SAMB) for the bond cluster [117], we find that in the D4h point group, only the irreducible representation B2g yields a non-zero toroidal quadrupole, which precisely corresponds to $T_{xy}$.
@@ Appendix C Intrinsic nonlinear current-induced spin and orbital magnetization
In this section, we provide a brief introduction to the formalism of the intrinsic nonlinear Edelstein effect. For a 2D system with either $C_{2z}$ or $\mathcal{P}$ symmetry, the linear spin and orbital responses are forbidden by symmetry. In such cases, the leading contribution arises at the second order:
@@
where $X_{a}$ denotes the $a$-th component of either the orbital angular momentum $L_{a}$ or the spin angular momentum $S_{a}$. The nonlinear response tensor $\chi_{abc}$ can be decomposed according to its parity under time-reversal symmetry [90, 91, 92, 125, 126]:
$\begin{split}\chi_{abc}^{\mathrm{even}}&=\tau\sum_{n}\int[d{\bm{k}}]f_{0}\partial_{c}\gamma^{X,n}_{ab},\\ \chi_{abc}^{\mathrm{odd}}&=\frac{1}{2}\int[d{\bm{k}}]\sum_{n}\Bigg[\Lambda_{abc,n}^{X}f_{0}-(X_{a,n}G_{bc,n}-v_{b,n}\mathscr{G}^{X}_{ac,n}-v_{c,n}\mathscr{G}^{X}_{ab,n})f^{\prime}_{0}\Bigg].\end{split}$ (11)
@@
The Zeeman-like field $\boldsymbol{h}$ couples to the spin or orbital degree of freedom in the form of $\boldsymbol{\hat{s}}\cdot\boldsymbol{h}$ or $\boldsymbol{\hat{L}}\cdot\boldsymbol{h}$. Therefore, in the presence of an external field, one can define the Berry connection and the related quantities in the $h$-space [90]. For example, the $h$-space Berry connection is defined as $\mathfrak{A}_{i}=\langle u_{n}({\bm{k}})|i\partial_{h_{i}}|u_{n}({\bm{k}})\rangle$. The Berry connection polarizability (BCP) and the $h$-space BCP are defined as
$\begin{split}G_{ab,n}&=2\textnormal{Re}\sum_{m\neq n}\frac{v_{a,nm}v_{b,mn}}{(\epsilon_{n}-\epsilon_{m})^{3}},\\ \mathscr{G}^{S}_{ab,n}&=-2\textnormal{Re}\sum_{m\neq n}\frac{s_{a,nm}v_{b,mn}}{(\epsilon_{n}-\epsilon_{m})^{3}},\\ \mathscr{G}^{O}_{ab,n}&=-2\textnormal{Re}\sum_{m\neq n}\frac{M_{a,nm}v_{b,mn}}{(\epsilon_{n}-\epsilon_{m})^{3}},\end{split}$ (12)
where the numerator involves the interband matrix elements of spin and velocity operators, $s_{a,nm}$ and $v_{a,nm}$, $M_{mn}=\sum_{l\neq n}(v_{ml}+\delta_{lm}v_{nn})\times A_{ln}/2$ is the interband orbital magnetic moment, $A_{mn}=\langle u_{m}|i\partial_{\bm{k}}|u_{n}\rangle$ is the Berry connection. The anomalous spin polarizability (ASP) and the anomalous orbital polarizability (AOP) is given by
$\begin{split}\gamma^{S,n}_{ab}({\bm{k}})&=2g\mu_{B}\textnormal{Im}\sum_{m\neq n}\frac{s_{a,nm}v_{b,mn}}{(\epsilon_{n}-\epsilon_{m})^{2}},\\ \gamma^{L,n}_{ab}({\bm{k}})&=-2\textnormal{Im}\sum_{m\neq n}\frac{M_{a,nm}v_{b,mn}}{(\epsilon_{n}-\epsilon_{m})^{2}}-\frac{e}{2\hbar}\epsilon_{akl}\Gamma_{n}^{lkb},\end{split}$ (13)
where the quantum Christoffel symbol is defined as
$\begin{split}\Gamma_{n}^{abc}&=\frac{1}{2}(\partial_{c}g_{n}^{ba}+\partial_{b}g_{n}^{ac}-\partial_{a}g_{n}^{bc}),\\ g_{n}^{ab}&=\mathrm{Re}\sum_{m\neq n}A_{nm}^{i}A_{mn}^{j}.\end{split}$ (14)
@@
$\begin{split}\langle u_{n}|\hat{\boldsymbol{L}}|u_{m}\rangle&=i\frac{e\hbar^{2}}{4g_{L}\mu_{B}}\sum_{q\neq n,m}(\frac{1}{\epsilon_{q}-\epsilon_{n}}+\frac{1}{\epsilon_{q}-\epsilon_{m}})\cdot\langle u_{n}|\hat{\boldsymbol{v}}|u_{q}\rangle\times\langle u_{q}|\hat{\boldsymbol{v}}|u_{m}\rangle.\end{split}$ (16)
@@
Note that the matrix elements of the orbital angular momentum (OAM) here are different from those of the atom-centered approximation (ACA), which assumes that the OAM originates from the region centered on the atomic sites. Within ACA, $\hat{\boldsymbol{L}}$ reduces to the on-site atomic orbital angular momentum operator that mixes orbitals located at the same lattice site [112, 127]. While the ACA is often adequate for wide-gap semiconductors, it is known to fail in both narrow-gap systems and in transition metals.